40 ft CL 5 ft A baseball pitching machine throws baseballs with a horizontal velocity V.. If you want the height h to be 42 in. (3.5 ft), determine the value of V.
Answer #1

Here, Time of travel t = sqrt(2s/g) = sqrt(2*(5-3.5)/32.2) = 0.3052 s

v0 = 40ft/0.3052s =   131 ft/s answer